# Magpie: Alignment Data Synthesis from Scratch by Prompting Aligned LLMs with Nothing

### v1

@@ E.2 Ablation Analysis on Filter Designs
Table 9: This table compares the performance of different filter designs within Magpie-Pro. All models are instruction-tuned on the Llama-8B base models.
@@
E.3 Performance of Magpie on More Benchmarks
We report the performance of models fine-tuned using Magpie-Air and Magpie-Pro, evaluated across a range of tasks featured on the Huggingface Open LLM Leaderboard [6] in Table 10. The tasks includes MMLU [22], ARC [11], HellaSwag [63], TruthfulQA [36], Winogard [30], and GSM8K [12]. We also perform experiments on MMLU-Redux [21] with zero-shot prompting. We use the default greedy decoding with $RP=1$ for all setups. Our experimental results demonstrate that models fine-tuned with Magpie-Air and Magpie-Pro achieve comparable performance to the official instruct model and other baselines despite the alignment tax.
Table 10: This table compares the performance of models instruction-tuned on Magpie-Air and Magpie-Pro against baselines and official instruct model across various downstream benchmarks. All models are instruction-tuned on the Llama-8B base models.
Alignment Setup MMLU (5) ARC (25) HellaSwag (10) TruthfulQA (0) Winograd (5) GSM8K (5) MMLU-Redux (0)
ShareGPT 66.03 58.45 81.50 52.34 74.03 48.67 50.68
Evol Instruct 65.62 60.75 82.70 52.87 76.16 42.91 52.73
OpenHermes 65.42 62.29 82.15 50.85 75.61 47.16 46.07
Tulu V2 Mix 66.34 59.22 82.80 47.99 76.16 58.07 46.97
WildChat 65.95 59.22 81.39 53.18 75.30 48.75 52.59
UltraChat 65.23 62.12 81.68 52.76 75.53 50.57 50.75
Magpie-Air-300K-Filtered 64.45 61.01 79.90 53.48 72.38 52.24 52.34
Magpie-Pro-100K-Filtered 65.31 60.32 81.18 51.11 73.32 50.42 52.56
Magpie-Pro-200K-Filtered 64.98 61.26 80.71 51.82 73.16 47.76 51.44
Magpie-Pro-300K-Filtered 64.25 60.41 80.52 52.46 73.32 47.92 52.16
Llama-3-8B-Instruct 67.82 61.52 78.67 52.47 72.14 71.72 58.60
Appendix F Prompt Templates
F.1 Prompt Templates for Magpie Extension
@@
Figure 11: Prompt for generating Magpie-MT. The placeholder {instruction} and {response} are from the first turn.
Prompt for controlling instruction tasks <|begin_of_text|><|start_header_id|>system<|end_header_id|> You are an AI assistant designed to provide helpful, step-by-step guidance on solving math problems. The user will ask you a wide range of complex mathematical questions. Your purpose is to assist users in understanding mathematical concepts, working through equations, and arriving at the correct solutions.<|eot_id|><|start_header_id|>user<|end_header_id|>
Figure 12: Prompt for controlling instruction tasks. In this example, we control LLMs to generate instructions related to math.
F.2 Prompt Templates for Evaluation
@@
Appendix G Magpie Examples
In this section, we demonstrate several representative instructions from Magpie-Pro with different task categories.

### v2

@@ F.4 Ablation Analysis on Filter Designs
Table 13: This table compares the performance of different filter designs within Magpie-Pro. All models are supervised-fine-tuned on the Llama-8B base models.
@@
F.5 Ablation Analysis on Response Generator
To investigate the impact of the response generator on the supervised fine-tuning performance using Magpie, we conduct an ablation study by replacing the response generator with Qwen-2-7B-Instruct (Yang et al., 2024) within Magpie-Air-300K-Filtered. We note that the performance of Qwen-2-7B-Instruct is comparable to, or slightly weaker than, Llama-3-8B-Instruct. The results are summarized in Table 14.
We observe that although there is a slight performance degradation, the model fine-tuned using Qwen-2-7B-Instruct as the response generator still outperforms all baselines, including those using GPT-4 as the response generator. These findings indicate two key points: (1) The success of Magpie depends little on the specific response generator used, and (2) the instructions generated by Magpie are of high quality and diversity.
Table 14: This table compares the impact of different response generators on the model performance. All models are supervised-fine-tuned on the Llama-8B base models.
Response Generator AlpacaEval 2 Arena-Hard
GPT-4-Turbo (1106) Llama-3-8B-Instruct
LC (%) WR (%) SD LC (%) WR (%) SD WR (%)
Llama-3-8B-Instruct 22.66 23.99 1.24 49.27 50.80 1.44 14.9
Qwen2-7B-Instruct 15.01 15.60 1.05 41.09 42.07 1.47 13.7
F.6 Trustworthiness of Magpie-Aligned Models
In what follows, we conduct more experiments to compare Magpie model and Llama-3-8B-Instruct on the TrustLLM benchmark (Huang et al., 2024). The results for safety, fairness, ethics, privacy, and robustness are summarized in Table 15.
We observe that our supervised-fine-tuned model slightly underperforms Llama-3-8B-Instruct in terms of safety and fairness. However, it outperforms the official instruct model on ethics, privacy, and robustness. Considering that our fine-tuned model uses much fewer data samples (300K compared to over 10M), these results again highlight the high quality of data generated by Magpie.
Table 15: This table compares the performance of model supervised-fine-tuned using Magpie-Pro-300K-Filtered and the official Llama-3-8B-Instruct on the TrustLLM benchmark (Huang et al., 2024).
TrustLLM Evaluation/Dataset Llama-3-8B-Instruct Magpie-Pro-300K-Filtered
[…]
Disparagement Sex ($\uparrow$) 0.99 0.99
Disparagement Race ($\uparrow$) 0.55 0.47
[…]
Appendix G Prompt Templates
G.1 Prompt Templates for Magpie Extension
@@
Figure 14: Prompt for generating Magpie-MT. We take Llama-3-8B-Instruct as an example. The placeholder {instruction} and {response} are from the first turn.
System Prompt Template <|begin_of_text|><|start_header_id|>system<|end_header_id|> {System Prompt}<|eot_id|><|start_header_id|>user<|end_header_id|>
System prompt for controlling math instruction tasks You are an AI assistant designed to provide helpful, step-by-step guidance on solving math problems. The user will ask you a wide range of complex mathematical questions. Your purpose is to assist users in understanding mathematical concepts, working through equations, and arriving at the correct solutions.
System prompt for controlling code instruction tasks You are an AI assistant designed to provide helpful, step-by-step guidance on coding problems. The user will ask you a wide range of coding questions. Your purpose is to assist users in understanding coding concepts, working through code, and arriving at the correct solutions.
System prompt for controlling translation tasks You are an AI assistant designed to provide accurate and contextually appropriate translations. Users will ask you to translate text between various languages. Your purpose is to assist users in understanding and conveying meaning across languages, maintaining the original context and nuances.
{CJK*}
UTF8gbsn
System prompt for controlling multilingual instruction generation (Japanese + Math) あなたはAIアシスタントで、数学の問題を解くために役立つ、ステップバイステップのガイダンスを提供するように設計されています。
Figure 15: Prompts for controlling instruction generation tasks. These examples illustrate how to guide Llama-3-8B-Instruct in generating instructions for specific domains: mathematics, coding, translation, and multilingual tasks. To adapt this approach for different instruction tasks, replace the System Prompt placeholder in the System Prompt Template with the appropriate domain-specific prompt.
G.2 Prompt Templates for Evaluation
@@
Appendix H Magpie Examples
In this section, we demonstrate several representative instructions generated by Magpie with different task categories. The following instructions are generated by the Llama-3-70B-Instruct model.
@@ Appendix H Magpie Examples
Magpie can also generate domain-specific instructions using models that are tailored to particular fields, as mentioned in Section 2.2. The following instructions are generated by DeepSeek-Coder-V2 (Zhu et al., 2024) and Qwen2-Math-7B-Instruct (Yang et al., 2024), respectively.
DeepSeek-Coder-V2 (Code Instruction) You are given a list of emails. You need to write a Python function that returns the domain, excluding the @ symbol, for each email.
Qwen2-Math-7B-Instruct (Math Instruction) A rectangle with length 12 units and width 8 units is scaled by a factor of 2 to form a new rectangle. Determine the dimensions of the new rectangle and calculate its area. Compare the area of the new rectangle to the area of the original rectangle.
We note that Magpie’s capabilities extend beyond generating English datasets to producing diverse multilingual datasets. The following instructions are generated by the Qwen2-72B-Instruct model.
{CJK*}
UTF8gbsn
Chinese 从给定的两个整数中找到较大的一个。但是你不能使用任何比较操作符(如>, <, !=等)或数学运算符(如+, -, *, /等)来实现它。你只能使用位操作符和逻辑操作符。
German Das Debate-System ’Oxford-Oberhaus’ wird bei ersten Auseinandersetzungen verwendet. Bitte erklären sie, wie dieses System funktioniert.
Spanish Según la encuesta anual de satisfacción al cliente que acabamos de realizar, parece que la satisfacción general de los clientes con nuestro rendimiento ha disminuido. ¿Podrías preparar una presentación detallada para la reunión del lunes que analice los resultados, identifique las áreas problemáticas y proporcione posibles soluciones basadas en los datos recogidos?
Portuguese Ho comprato una nuova inchiostriera sulla quale è presente la scritta "Non manipolare". Cosa signidica?
Italian Crie um exemplo de uma conversa entre dois personagens, um MC de hip hop e um pianista clássico, discutindo sobre seus estilos favoritos de música.